# Memory Decoder: A Pretrained, Plug-and-Play Memory for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have shown strong abilities in general language tasks, yet adapting them to specific domains remains a challenge. Current method like Domain Adaptive Pretraining (DAPT) requires costly full-parameter training and suffers from catastrophic forgetting. Meanwhile, Retrieval-Augmented Generation (RAG) introduces substantial inference latency due to expensive nearest-neighbor searches and longer context. This paper introduces Memory Decoder, a plug-and-play pretrained memory that enables efficient domain adaptation without changing the original model’s parameters. Memory Decoder employs a small transformer decoder that learns to imitate the behavior of an external non-parametric retriever. Once trained, Memory Decoder can be seamlessly integrated with any pretrained language model that shares the same tokenizer, requiring no model-specific modifications. Experimental results demonstrate that Memory Decoder enables effective adaptation of various Qwen and Llama models to three distinct specialized domains: biomedicine, finance, and law, reducing perplexity by an average of 6.17 points. Overall, Memory Decoder introduces a novel paradigm centered on a specially pretrained memory component designed for domain-specific adaptation. This memory architecture can be integrated in a plug-and-play manner, consistently enhancing performance across multiple models within the target domain.  [^1] Code and checkpoints will be released upon acceptance.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks (Grattafiori et al., 2024; Yang et al., 2024; Liu et al., 2024; Guo et al., 2025). Pretrained on vast corpora of general text data, LLMs have revolutionized how we approach language understanding and generation tasks. However, despite their impressive general capabilities, adapting LLMs to perform optimally in specific domains remains a significant challenge. Domain-specific adaptation is crucial for applications in specialized fields such as biomedicine, finance, and law (Chen et al., 2023; Liu et al., 2023b; Colombo et al., 2024), where domain expertise and terminology are essential for accurate and reliable performance.
@@
Domain adaptation for pretrained language models has traditionally followed several approaches, each with distinct advantages and limitations. Domain Adaptive Pre-Training (DAPT) involves continued pre-training of the LLM on domain-specific corpora (Gururangan et al., 2020). While effective, this approach suffers from substantial computational costs associated with full-parameter training, especially as model sizes continue to grow into billions of parameters. Furthermore, adapting multiple models to the same domain requires separate training runs for each model, leading to resource inefficiency. Even with successful DAPT implementation, these models often encounter catastrophic forgetting, where the adaptation process diminishes the model’s general capabilities (Kirkpatrick et al., 2017; Ven van de et al., 2024).
Retrieval-Augmented Generation (RAG) offers an alternative approach by enhancing model outputs with relevant retrieved information (Lewis et al., 2020a; Izacard et al., 2023a). While this method preserves the original model parameters, it introduces substantial computation overhead during inference due to expensive nearest neighbor (kNN) searches across large datastores and extended context (He et al., 2021a).
These two approaches present a fundamental dilemma in domain adaptation: DAPT requires costly training procedures and cannot efficiently adapt multiple models to the same domain, while RAG introduces significant computation and storage overhead during inference. This inherent trade-off between the plug-and-play nature of RAG and the inference efficiency of DAPT highlights the research gap for a solution offering both adaptability across models and computational efficiency during deployment. To address this challenge, we propose Memory Decoder (MemDec), a plug-and-play pretrained memory designed for efficient domain adaptation of large language models without modifying their parameters. Our approach draws inspiration from retrieval-based methods like kNN-LM (Khandelwal et al., 2019a), but overcomes their limitations through a different paradigm. Rather than building and searching model-specific datastores during inference, Memory Decoder employs a small transformer decoder that is specially pretrained to imitate the behavior of non-parametric retrievers by aligning its output distribution with the ones of non-parametric retrievers. Figure 1 illustrates how our approach differs fundamentally from both DAPT and RAG.
@@ 2.2 Nearest Neighbor Language Models
The $k$-nearest neighbor language model (kNN-LM) (Khandelwal et al., 2019a) enables non-parametric domain adaptation without modifying the pretrained model’s parameters.
@@
While effective, kNN-LM introduces substantial computational and storage overhead during inference. For instance, the Wikitext-103 datastore requires nearly 500GB storage even for GPT2-small model (He et al., 2021a). These limitations motivate our Memory Decoder, a compact parametric model pretrained to mimic retrieval behavior while eliminating the need for large datastores.
@@ 3.1 Pre-training
Since we require non-parametric distributions as supervision signals, we construct training pairs of $(x_{i},p_{\text{kNN}}(\cdot|x_{i}))$ in advance to enable efficient pre-training. Here, $x_{i}$ represents the input context and $p_{\text{kNN}}(\cdot|x_{i})$ denotes the probability distribution generated by the non-parametric retriever for that context. First, we build a key-value datastore $(K,V)=\{(\phi(x_{i}),y_{i})\mid(x_{i},y_{i})\in\mathcal{D}_{\text{train}}\}$ using our domain-specific corpus, where $\phi(\cdot)$ extracts hidden representations from a specific layer of the pretrained model. For each context $x_{i}$ in the corpus, we then perform $k$-nearest neighbor search against this datastore to identify similar contexts. To avoid trivial self-retrieval that would contaminate the learning signal, we exclude the top-1 neighbor when its key exactly matches the query key. Finally, we compute the non-parametric distribution $p_{\text{kNN}}(\cdot|x_{i})$ for each context using the retrieved neighbors and cache these context-distribution pairs for training.
@@
Unlike traditional language modeling with single-label targets, kNN distributions offer richer supervision signals by capturing the diversity of plausible continuations in the domain (Xu et al., 2023)(see Appendix G for detailed analysis on kNN distributions). Through extensive experimentation, we have identified that a hybrid objective yields optimal performance.
Our approach centers on a Distribution Alignment Loss that minimizes the KL divergence (Van Erven, Harremos, 2014) between Memory Decoder’s output distribution and the cached kNN distributions for each sample:
@@
To prevent excessive deviation from the underlying corpus distribution, we integrate a complementary standard Language Modeling objective (Zhang, Sabuncu, 2018):
@@
Previous failed attempts to learn kNN distributions and our conjecture on why vanilla KL divergence with cross-entropy regularization succeeds are detailed in Appendix H.
@@ 3.2 Inference
Unlike traditional retrieval-augmented approaches that introduce substantial latency from nearest neighbor search and extended context processing, Memory Decoder requires only a single forward pass through a relatively small transformer decoder. As demonstrated in Figure 4, our method achieves significant improvements in inference efficiency compared to alternative domain adaptation techniques. With just 1.28× overhead relative to the base model, Memory Decoder substantially outperforms both In-Context RAG (Ram et al., 2023) (1.51×) and kNN-LM (Khandelwal et al., 2019a) (2.17×), with the efficiency gap widening as the number of processed tokens increases. Our method achieves even higher relative speedup (10.0×) when augmenting larger base models with a 500-million-entry datastore, as the inter-process communication overhead between Memory Decoder and LLM becomes amortized by longer inference times, unlike kNN search which scales linearly with datastore size. This computational advantage, combined with Memory Decoder’s model-agnostic design, makes our approach particularly valuable for production environments where both performance and efficiency are critical considerations.
Figure 4: Inference latency comparison across domain adaptation methods. These measurements were conducted on Qwen2.5-1.5B (Yang et al., 2024) for biomedicine domain text, augmented by a 0.5B Memory Decoder.
@@ 4 Experimental Setup
We evaluate Memory Decoder across six complementary settings: (1) Language modeling on WikiText-103 (§5.1) to demonstrate effectiveness across GPT-2 model scales; (2) Downstream tasks (§5.2) to verify preservation of general capabilities during domain adaptation; (3) Cross-model adaptation (§5.3) showing a single Memory Decoder enhancing Qwen models from 0.5B to 72B parameters; (4) Cross-vocabulary adaptation (§5.4) demonstrating efficient transfer between tokenizer families; and (5) Knowledge-intensive QA (§5.5) proving Memory Decoder maintains reasoning abilities while enhancing factual recall—a key limitation of retrieval methods. (6) Domain-specific downstream tasks (Appendix D) confirming preservation of in-context learning capabilities across 13 real-world benchmarks. These experiments establish Memory Decoder as a versatile, plug-and-play solution for efficient domain adaptation across diverse architectures and applications.
@@
For language modeling experiments, we use Wikitext-103 (Merity et al., 2016), a standard benchmark containing over 100M tokens from Wikipedia articles. For downstream evaluation, following the kNN-prompt framework, we assess performance across nine NLP tasks: sentiment analysis (SST2 (Socher et al., 2013), MR (Pang, Lee, 2005b), CR (Hu, Liu, 2004), RT (Pang, Lee, 2005a)), textual entailment (HYP (Kiesel et al., 2019), CB (De Marneffe et al., 2019), RTE (Dagan et al., 2010)), and text classification (AGN (Zhang et al., 2015a), Yahoo (Zhang et al., 2015b)). For domain-specific adaptation, we utilize three specialized corpora: (1) biomedical text from MIMIC-III (Johnson et al., 2016) clinical notes covering over 46,000 patients, (2) financial news (Liu et al., 2023a) from April 2024 to February 2025 for over 5,000 stocks, and (3) legal text from the Asylex corpus (Barale et al., 2023) containing 59,112 documents of refugee status determination in Canada from 1996 to 2022.
@@
We compare Memory Decoder against several established domain adaptation methods: In-Context RAG (Ram et al., 2023), which implements a BM25 retriever that processes 32 query tokens, with retrieval occurring every 4 tokens. kNN-LM (Khandelwal et al., 2019a), configured with interpolation parameter $\lambda=0.25$ and temperature settings of $\tau=1$ for GPT-2 small and medium, and $\tau=13$ for large and xl models. LoRA (Hu et al., 2022), applied to query, key, value and MLP layers, with rank adjusted for each model to achieve parameter counts comparable to Memory Decoder. Domain Adaptive Pretraining(DAPT) (Gururangan et al., 2020), which involves complete retraining of all model parameters on the domain-specific corpus.
@@
We conduct our experiments on an 8×A800 80GB GPU setup. For language modeling and downstream evaluations, we use GPT2-xl (Radford et al., 2019) to build the key-value datastore and non-parametric distributions for training, and continue training on a GPT2-small model with learning rate 1e-3. For cross-model adaptation, we use Qwen2.5-1.5B (Yang et al., 2024) to build the datastore, and continue training on Qwen2.5-0.5B with learning rate 1e-4. For cross-vocabulary adaptation, we use Llama3.2-1B (Grattafiori et al., 2024) to build the datastore, and continue training on the Memory Decoder trained from cross-model experiments, with its embedding layer and language model head re-initialized. All experiments use a training budget equivalent to the computational cost of training a 7B parameter model for 1 epoch, with DAPT and LoRA baselines using the same maximum training FLOPS but early stopped to prevent overfitting. The training hyperparameter $\beta$ is set to $0.5$ across all tasks.
@@
For language modeling, cross-model, and cross-tokenizer experiments, we use sliding window perplexity. Following Baevski, Auli (2018), in each test example, the context length is set to 1024 where only the latter 512 tokens are scored. For downstream evaluation, following methodology from Shi et al. (2022), we report results using the domain-conditional PMI scoring rule (Holtzman et al., 2021). The interpolation hyperparameter $\alpha$ is tuned on the validation split of each task following Khandelwal et al. (2019a), see more details in Appendix A.
@@ 5.1 Language Modeling on Wikitext-103
Table 1 demonstrates the exceptional effectiveness of Memory Decoder across all GPT2 model sizes. A single Memory Decoder with only 124M parameters consistently enhances the entire GPT2 family, showcasing its plug-and-play capability regardless of base model size. For smaller models, our approach delivers superior results compared to all adaptation methods—notably outperforming DAPT by 15.1% for GPT2-small and maintaining an advantage for GPT2-medium despite utilizing only a fraction of the parameters. Even when applied to larger models where DAPT has inherent advantages due to full model updates, Memory Decoder remains highly competitive while consistently outperforming all other parameter-efficient methods without modifying any original parameters. These results validate that a small parametric decoder can effectively capture the benefits of non-parametric retrieval while eliminating computational overhead.
@@ 5.2 Downstream Performance
Table 2 reveals Memory Decoder’s ability to enhance domain adaptation while preserving general language capabilities in zero-shot evaluation settings. Unlike DAPT, which suffers catastrophic forgetting on several tasks (particularly HYP and Yahoo where performance drops by nearly half; see Appendix B for detailed analysis), Memory Decoder maintains or improves performance across all evaluated tasks. Our approach achieves the highest average score across all nine tasks, outperforming the base model, kNN-LM, and LoRA while demonstrating particular strength on textual entailment tasks like CB and RTE. These results validate a key advantage of our plug-and-play architecture: by keeping the original model parameters intact while augmenting them with domain knowledge, Memory Decoder achieves domain adaptation without sacrificing general capabilities. Importantly, all experiments are conducted in a zero-shot setting, and our method should be viewed as orthogonal to in-context learning approaches, which we analyze in Appendix F.

### v2

@@ Abstract
Large Language Models (LLMs) have shown strong abilities in general language tasks, yet adapting them to specific domains remains a challenge. Current method like Domain Adaptive Pretraining (DAPT) requires costly full-parameter training and suffers from catastrophic forgetting. Meanwhile, Retrieval-Augmented Generation (RAG) introduces substantial inference latency due to expensive nearest-neighbor searches and longer context. This paper introduces Memory Decoder, a plug-and-play pretrained memory that enables efficient domain adaptation without changing the original model’s parameters. Memory Decoder employs a small transformer decoder that learns to imitate the behavior of an external non-parametric retriever. Once trained, Memory Decoder can be seamlessly integrated with any pretrained language model that shares the same tokenizer, requiring no model-specific modifications. Experimental results demonstrate that Memory Decoder enables effective adaptation of various Qwen and Llama models to three distinct specialized domains: biomedicine, finance, and law, reducing perplexity by an average of 6.17 points. Overall, Memory Decoder introduces a novel paradigm centered on a specially pretrained memory component designed for domain-specific adaptation. This memory architecture can be integrated in a plug-and-play manner, consistently enhancing performance across multiple models within the target domain.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks [Grattafiori et al., 2024, Yang et al., 2024, Liu et al., 2024, Guo et al., 2025]. Pretrained on vast corpora of general text data, LLMs have revolutionized how we approach language understanding and generation tasks. However, despite their impressive general capabilities, adapting LLMs to perform optimally in specific domains remains a significant challenge. Domain-specific adaptation is crucial for applications in specialized fields such as biomedicine, finance, and law [Chen et al., 2023, Liu et al., 2023b, Colombo et al., 2024], where domain expertise and terminology are essential for accurate and reliable performance.
Domain adaptation for pretrained language models has traditionally followed several approaches, each with distinct advantages and limitations. Domain Adaptive Pre-Training (DAPT) involves continued pre-training of the LLM on domain-specific corpora [Gururangan et al., 2020]. While effective, this approach suffers from substantial computational costs associated with full-parameter training, especially as model sizes continue to grow into billions of parameters. Furthermore, adapting multiple models to the same domain requires separate training runs for each model, leading to resource inefficiency. Even with successful DAPT implementation, these models often encounter catastrophic forgetting, where the adaptation process diminishes the model’s general capabilities [Kirkpatrick et al., 2017, van de Ven et al., 2024].
Retrieval-Augmented Generation (RAG) offers an alternative approach by enhancing model outputs with relevant retrieved information [Lewis et al., 2020, Izacard et al., 2023]. While this method preserves the original model parameters, it introduces substantial computation overhead during inference due to expensive nearest neighbor (kNN) searches across large datastores and extended context [He et al., 2021].
These two approaches present a fundamental dilemma in domain adaptation: DAPT requires costly training procedures and cannot efficiently adapt multiple models to the same domain, while RAG introduces significant computation and storage overhead during inference. This inherent trade-off between the plug-and-play nature of RAG and the inference efficiency of DAPT highlights the research gap for a solution offering both adaptability across models and computational efficiency during deployment. To address this challenge, we propose Memory Decoder (MemDec), a plug-and-play pretrained memory designed for efficient domain adaptation of large language models without modifying their parameters. Our approach draws inspiration from retrieval-based methods like kNN-LM [Khandelwal et al., 2019], but overcomes their limitations through a different paradigm. Rather than building and searching model-specific datastores during inference, Memory Decoder employs a small transformer decoder that is specially pretrained to imitate the behavior of non-parametric retrievers by aligning its output distribution with the ones of non-parametric retrievers. Figure 1 illustrates how our approach differs fundamentally from both DAPT and RAG.
@@ 2.2 Nearest Neighbor Language Models
The $k$-nearest neighbor language model (kNN-LM) [Khandelwal et al., 2019] enables non-parametric domain adaptation without modifying the pretrained model’s parameters.
@@
While effective, kNN-LM introduces substantial computational and storage overhead during inference. For instance, the Wikitext-103 datastore requires nearly 500GB storage even for GPT2-small model [He et al., 2021]. These limitations motivate our Memory Decoder, a compact parametric model pretrained to mimic retrieval behavior while eliminating the need for large datastores.
@@ 3.1 Pre-training
Since we require non-parametric distributions as supervision signals, we construct training pairs of $(x_{i},p_{\text{kNN}}(\cdot|x_{i}))$ in advance to enable efficient pre-training. Here, $x_{i}$ represents the input context and $p_{\text{kNN}}(\cdot|x_{i})$ denotes the probability distribution generated by the non-parametric retriever for that context. First, we build a key-value datastore $(K,V)=\{(\phi(x_{i}),y_{i})\mid(x_{i},y_{i})\in\mathcal{D}_{\text{train}}\}$ using our domain-specific corpus, where $\phi(\cdot)$ extracts hidden representations from a specific layer of the pretrained model. For each context $x_{i}$ in the corpus, we then perform $k$-nearest neighbor search against this datastore to identify similar contexts. To avoid trivial self-retrieval that would contaminate the learning signal, we exclude the top-1 neighbor where its key exactly matches the query key. Finally, we compute the non-parametric distribution $p_{\text{kNN}}(\cdot|x_{i})$ for each context using the retrieved neighbors and cache these context-distribution pairs for training.
@@
Unlike traditional language modeling with single-label targets, kNN distributions offer richer supervision signals by capturing the diversity of plausible continuations in the domain [Xu et al., 2023](see Appendix C for detailed analysis on kNN distributions). Through extensive experimentation, we have identified that a hybrid objective yields optimal performance.
Our approach centers on a Distribution Alignment Loss that minimizes the KL divergence [Van Erven and Harremos, 2014] between Memory Decoder’s output distribution and the cached kNN distributions for each sample:
@@
To prevent excessive deviation from the underlying corpus distribution, we integrate a complementary standard Language Modeling objective [Zhang and Sabuncu, 2018]:
@@
Previous failed attempts to learn kNN distributions and our conjecture on why vanilla KL divergence with cross-entropy regularization succeeds are detailed in Appendix D.
@@ 3.2 Inference
Unlike traditional retrieval-augmented approaches that introduce substantial latency from nearest neighbor search and extended context processing, Memory Decoder requires only a single forward pass through a relatively small transformer decoder. As demonstrated in Figure 4, our method achieves significant improvements in inference efficiency compared to alternative domain adaptation techniques. With just 1.28× overhead relative to the base model, Memory Decoder substantially outperforms both In-Context RAG [Ram et al., 2023] (1.51×) and kNN-LM [Khandelwal et al., 2019] (2.17×). This computational advantage, combined with Memory Decoder’s model-agnostic design, makes our approach particularly valuable for production environments where both performance and efficiency are critical considerations.
Figure 4: Inference latency comparison across domain adaptation methods. These measurements were conducted on Qwen2.5-1.5B [Yang et al., 2024] for biomedicine domain text, augmented by a 0.5B Memory Decoder.
@@ 4 Experimental Setup
We evaluate Memory Decoder across four complementary settings: (1) Language modeling on WikiText-103 (§5.1) to demonstrate effectiveness across GPT-2 model scales; (2) Downstream tasks (§5.2) to verify preservation of general capabilities during domain adaptation; (3) Cross-model adaptation (§5.3) showing a single Memory Decoder enhancing Qwen models from 0.5B to 72B parameters; (4) Cross-vocabulary adaptation (§5.4) demonstrating efficient transfer between tokenizer families; These experiments establish Memory Decoder as a versatile, plug-and-play solution for efficient domain adaptation across diverse architectures and applications.
@@
For language modeling experiments, we use Wikitext-103 [Merity et al., 2016], a standard benchmark containing over 100M tokens from Wikipedia articles. For downstream evaluation, following the kNN-prompt framework, we assess performance across nine NLP tasks: sentiment analysis (SST2 [Socher et al., 2013], MR [Pang and Lee, 2005a], CR [Hu and Liu, 2004], RT [Pang and Lee, 2005b]), textual entailment (HYP [Kiesel et al., 2019], CB [De Marneffe et al., 2019], RTE [Dagan et al., 2010]), and text classification (AGN [Zhang et al., 2015a], Yahoo [Zhang et al., 2015b]). For domain-specific adaptation, we utilize three specialized corpora: (1) biomedical text from MIMIC-III [Johnson et al., 2016] clinical notes covering over 46,000 patients, (2) financial news [Liu et al., 2023a] from April 2024 to October 2024, and (3) legal text from the Asylex corpus [Barale et al., 2023] containing 59,112 documents of refugee status determination in Canada from 1996 to 2022.
@@
We compare Memory Decoder against several established domain adaptation methods: In-Context RAG [Ram et al., 2023], which implements a BM25 retriever that processes 32 query tokens, with retrieval occurring every 4 tokens. kNN-LM [Khandelwal et al., 2019], configured with interpolation parameter $\lambda=0.25$ and temperature settings of $\tau=1$ for GPT-2 small and medium, and $\tau=13$ for large and xl models. LoRA [Hu et al., 2022], applied to query, key, value and MLP layers, with rank adjusted for each model to achieve parameter counts comparable to Memory Decoder. Domain Adaptive Pretraining(DAPT) [Gururangan et al., 2020], which involves complete retraining of all model parameters on the domain-specific corpus.
@@
We conduct our experiments on an 8×A800 80GB GPU setup. For language modeling and downstream evaluations, we use a GPT2-xl model(finetuned on wikitext) to build the key-value datastore and non-parametric distributions for training, and continue training on a GPT2-small model(finetuned on wikitext) with learning rate 1e-3. For cross-model adaptation, we use Qwen2.5-1.5B [Yang et al., 2024] to build the datastore, and continue training on Qwen2.5-0.5B with learning rate 1e-4. For cross-vocabulary adaptation, we use Llama3.2-1B [Grattafiori et al., 2024] to build the datastore, and continue training on the Memory Decoder trained from cross-model experiments, with its embedding layer and language model head re-initialized. All experiments use a training budget equivalent to the computational cost of training a 7B parameter model for 1 epoch, with DAPT and LoRA baselines using the same maximum training FLOPS but early stopped to prevent overfitting. The training hyperparameter $\beta$ is set to $0.5$ across all tasks.
@@
For language modeling, cross-model, and cross-tokenizer experiments, we use sliding window perplexity. Following Baevski and Auli [2018], in each test example, the context length is set to 1024 where only the latter 512 tokens are scored. For downstream evaluation, following methodology from Shi et al. [2022], we report results using the domain-conditional PMI scoring rule [Holtzman et al., 2021]. The interpolation hyperparameter $\alpha$ is tuned on the validation split of each task following Khandelwal et al. [2019], see more details in Appendix A.
@@ 5.1 Language Modeling on Wikitext-103
Table 1 demonstrates the exceptional effectiveness of Memory Decoder across all GPT2 model sizes. A single Memory Decoder with only 124M parameters consistently enhances the entire GPT2 family, showcasing its plug-and-play capability regardless of base model size. For smaller models, our approach delivers superior results compared to all adaptation methods—notably maintaining an advantage for GPT2-medium despite utilizing only one third of the parameters. Even when applied to larger models where DAPT has inherent advantages due to full model updates, Memory Decoder remains highly competitive while consistently outperforming all other parameter-efficient methods without modifying any original parameters. These results validate that a small parametric decoder can effectively capture the benefits of non-parametric retrieval while eliminating computational overhead.
@@ 5.2 Downstream Performance
Table 2 reveals Memory Decoder’s ability to enhance domain adaptation while preserving general language capabilities in zero-shot evaluation settings. Unlike DAPT, which suffers catastrophic forgetting on several tasks (particularly HYP and Yahoo where performance drops by nearly half; see Appendix B for detailed analysis), Memory Decoder maintains or improves performance across all evaluated tasks. Our approach achieves the highest average score across all nine tasks, outperforming the base model, kNN-LM, and LoRA while demonstrating particular strength on textual entailment tasks like CB and RTE. These results validate a key advantage of our plug-and-play architecture: by keeping the original model parameters intact while augmenting them with domain knowledge, Memory Decoder achieves domain adaptation without sacrificing general capabilities. Importantly, all experiments are conducted in a zero-shot setting, and our method should be viewed as orthogonal to in-context learning approaches.